# Multi-UAV Behavior-based Formation with Static and Dynamic Obstacles Avoidance via Reinforcement Learning

### v1

@@ V-A2 The Second Stage, Reward Generalization
In the first stage, we find a proper w in a simplified setting. Now, we use the weight to scalarize the reward vector and tackle a more complex task: three drones navigating through both static and dynamic obstacles. This stage confirms the generalizability of the derived reward function and highlights the effectiveness of our proposed approach.
@@
Training the RL policy from scratch for complex tasks is computationally intensive and often leads to suboptimal solutions. Hence, we utilize curriculum learning to speed up the training process through gradually increasing task difficulty. Specifically, we implement a 3-period curriculum, where drones first learn to fly forward in an obstacle-free environment, followed by a static-obstacle-only environment, and ultimately a mixed-obstacle environment.
V-B Network Architecture
V-B1 Attention-Based Observation Encoder
Figure 3: The network structure of the observation encoder.
@@ V-B1 Attention-Based Observation Encoder
To efficiently manage scenarios with different quantities of obstacles, we develop an attention-based observation encoder. This encoder eliminates the limitations on input dimensionality and emphasizes relevant task information. The structure of the encoder is shown in Fig. 3. The observation composes of four parts: $\mathbf{o}_{self}$, $\mathbf{o}_{drones}$, $\mathbf{o}_{static}$, and $\mathbf{o}_{dynamic}$. Each part is individually encoded through separate MLPs to produce embeddings of identical dimensions. We use multi-head self-attention on these embeddings to obtain respective features. To better capture the relationships between the drone and other environmental entities, the features are processed through a multi-head cross-attention module, utilizing self-feature as the query and other features as the key and value. The final feature $\mathbf{f}$ is a concatenation of the self-feature and the output of the multi-head cross-attention.
V-B2 Actor&Critic Network
@@
Following the observation encoder, a Gaussian distribution model $\pi_{\theta}=\mathcal{N}(\bm{\mu}_{\theta},\bm{\sigma}_{\theta})$ is used to parameterize actions in the actor network. Additionally, an MLP outputs a scalar value $V=V_{\phi}(\mathbf{f})$ as the estimated reward-to-go for the current state in the critic network.
@@ VI Experiments
VI-A Baselines
Method Type Directed Flight Formation Maintenance Obstacle Avoidance Action Type
Swarm-Formation [37] Planning Static Trajectory
R-Mader [30] Planning Static & Dynamic (Trajectory Needed) Trajectory
Swarm-RL [34] RL Static Thrust
Ours RL Static & Dynamic CTBR
TABLE I: Characteristics of baselines and our method.
We compare our approach with three state-of-the-art baselines that meet some objectives: Swarm-Formation [37], R-Mader [30], and Swarm-RL [34]. We outline the characteristics of the baselines in Tab. I and leave the modifications that tailor them to our task to the website. Since R-Mader and Swarm-RL do not inherently manage formation, we set the desired formation as the target positions of the drones to encourage formation maintenance.
VI-B Simulation Results
We choose OmniDrones [38] as our testbed, a drone simulator designed for RL policy training, thanks to its high-speed GPU simulation and complex environment dynamics for drones.
We consider a challenging task: 3 drones fly as a near-equilateral triangle while avoiding static and dynamic obstacles scattered within the space. We obtain the policy through our proposed 2-stage RL training pipeline. During the first training stage, each trial requires 100 million environment steps. During the second training stage, we perform 3-period curriculum learning, where we train the RL policy in scenarios with no obstacle, 10 static obstacles only, and mixed obstacles with 10 static and 2 dynamic obstacles, sequentially. Each period takes 15 million, 150 million, and 150 million environment steps, respectively. Swarm-RL takes 1 billion environment steps to converge, about 3x steps of our method.
VI-B1 Evaluation Metrics
@@
  • Formation Maintenance (FM). In successful episodes, we compute the unnormalized Laplacian distance between the target formation and the actual swarm configuration. A smaller unnormalized Laplacian distance indicates a more desirable formation in terms of both shape and size.
Note that we do not quantify the performance of action smoothing, since this term facilitates sim2real transfer, which will be verified in real-world experiments. In the following section, we report the CFR and FM of planning-based methods averaged over 25 repeated experiments, as they do not support parallel evaluation, and RL-based methods over 100 trials. The positions of all obstacles are randomly generated in each trial.
@@
TABLE II: Performance of all methods under mixed obstacle scenario (2 balls + 10 columns) in simulation.
@@
VI-B2 Main Results
We test all methods in the mixed obstacle scenario with 2 balls and 10 columns and report the performance in Tab. II. We further test the zero-shot transfer ability of our policy by varying the number of obstacles. We gradually increase the number of columns from 5 to 20 under static obstacle scenario, and the number of balls from 1 to 5 under dynamic obstacle scenario. The results are shown in Fig. 4 and Fig. 5, respectively.
Our method outperforms all baseline methods regarding both CFR and FM in the mixed obstacle scenario, and demonstrates comparable, if not superior, performance to the best baseline in extremely confined environments with 20 columns or 5 balls. The results demonstrate our method’s strong capability to maintain effective formation while evading various obstacles, underscoring its ability to accomplish multi-objective tasks. Also, as the number of obstacles increases, our method shows only a slight decrease in both metrics, demonstrating its strong scalability and generalization capability, which we attribute to the attention mechanism employed in the network structure.
R-Mader, which inherently incorporates dynamic obstacle avoidance, performs better in dynamic scenarios compared to Swarm-Formation, but still falls short of our method’s performance. This is because drones can only observe obstacles within a certain distance in our task setting. For R-Mader, this is equivalent to reducing the response time for obstacle avoidance.
VI-B3 Behavior Visualization
In Fig. 6, we visualize the formation behavior of our method under static scenario with 4 columns (represented by black boxes) and dynamic scenario with 1 ball (represented by red circles). In static scenario, our method adjusts the centroid of the formation and slightly twists its formation, demonstrating the flexibility of behavior-based formation structure. In dynamic scenario, our method requires all drones to decelerate upon encountering the ball and resume flight after the ball drops.
(a) Static obstacle scenario.
(b) Dynamic obstacle scenario.
Figure 6: Behavior visualization. (Top view)
VI-C Analysis of Reward Scalarization
Fig. 7 (a) illustrates the satisfaction rates (SR) of policies trained with randomly sampled weight vectors w. We observe that weight vectors with similar SR can yield diverse behaviors. Fig. 7 (b) depicts the trajectories of 3 drones dodging a dynamic obstacle, guided by policies with similar satisfaction rates. When encountering a dynamic obstacle, the first policy and the third policy decelerate the drones in different magnitudes, whereas the second policy chooses to ascend and accelerate the drones. This suggests that our method holds strong flexibility and policy diversity, allowing humans to choose behaviors that best align with their preferences. In practice, we select the weight vector with the highest SR.
@@ VI-C Analysis of Reward Scalarization
VI-D Ablation Studies
@@
In this section, we conduct ablation studies on the attention-based observation encoder and CL used in the second stage. All experiments are conducted in mixed obstacle scenario with 10 columns and 2 balls. We set success criteria for each objective and record the proportion of episodes meeting these criteria. Furthermore, we use Success Rate to record the proportion of episodes meeting all criteria simultaneously.
VI-D1 Attention-based Observation Encoder
We compare three variants: (1) Ours w/o CA, where the cross-attention module is replaced with an MLP, (2) Ours w/o SA, where the self-attention module is replaced with an MLP, and (3) MLP, where both the cross-attention and self-attention modules are substituted with an MLP.
@@ VI-D2 Curriculum Learning
VI-E Real World Deployment
We deploy our RL policy and the planning-based baselines on Crazyflie 2.0. Due to the limited processing power of Crazyflie’s onboard processors, we execute the RL policy and planning baselines on a computer, and send commands to the drones via Crazyradio. We acquire the ground truth positions of the drones and obstacles through a motion capture system and estimate their linear and angular velocities by calculating the differences in position and rotation between consecutive frames. For the RL policy, CTBR commands are transmitted via radio at 50 Hz, while for planning baselines, trajectory waypoints are sent through position commands at 10 Hz. We tried to deploy Swarm-RL both on-board and on external PC, but failed in both cases; therefore, we do not present the real-world results of Swarm-RL.
VI-E1 Sim2Real Transfer
For better Sim2Real transfer, we add domain randomization and CTBR clip during training. More specifically, we randomize the initial position of each drones within 0.05m of the original formation, and randomize the initial rotation within $[-0.1\pi,0.1\pi]$. Also, to avoid extreme actions, we clip the policy output, setting $\omega$ within $[-45,45]$ degree/s and $c$ within $[0.4,0.9]$.
VI-E2 Results and Analysis
In real world deployment, we consider three scenarios, i.e., with 2 static columns, with a manually thrown ball, and with both the columns and the ball. For full video, please refer to the supplementary materials.
The results further demonstrate the strong Sim2Real transfer capability of our method. Our method can perform aggressive dodging strategy to safely navigate through all obstacles while maintaining formation, whereas the baseline methods fail to find a viable path. Swarm-Formation halts when the dynamic obstacle drops before the drone; R-Mader struggles with cluttered static obstacles and fails to keep formation. Through the zero-shot Sim2Real deployment, we validate the efficacy of the action smoothing objective.
@@ VII Conclusion and Future Work
In this paper, we employ multi-agent reinforcement learning to tackle the challenging task of maintaining behavior-based formation among multiple UAVs while avoiding both static and dynamic obstacles during directed flight. Our approach involves a two-stage training pipeline. In the first stage, we search for a balanced weight of all objectives in a simplified task setting. In the second stage, we apply curriculum learning to navigate the large exploration space, deriving a feasible policy for the more complex task. We introduce an attention-based observation encoder to effectively capture the features of the drones and the environments, ensuring the scalability of the policy. Our approach outperforms SOTA baselines in terms of both CFR and FM. The effectiveness of our method is further validated through real-world deployment.

### v2

@@ V-B2 The Second Stage, Reward Generalization
With a w that properly scalarizes R, we can now tackle a more complex task: three drones navigating through both static and dynamic obstacles. This stage demonstrates the generalizability of the derived reward function and highlights the effectiveness of our proposed approach.
@@
Training the RL policy from scratch for complex tasks is computationally intensive and often leads to suboptimal solutions. Hence, we utilize curriculum learning to speed up training through gradually increasing task difficulty. Specifically, we implement a 3-period curriculum, where drones first learn to fly forward in an obstacle-free environment, followed by a static-obstacle-only environment, and ultimately a mixed-obstacle environment.
V-C Network Architecture
V-C1 Attention-Based Observation Encoder
Figure 2: The network structure of the observation encoder. The subscripts sf, dr, st, and dy are abbreviations for self, drones, static, and dynamic, as mentioned in the main text.
@@ V-C1 Attention-Based Observation Encoder
To efficiently manage scenarios with different quantities of obstacles, we develop an attention-based observation encoder. This encoder eliminates the limitations on input dimensionality and emphasizes relevant task information. The structure of the encoder is shown in Fig. 2. The observation composes of four parts: $\mathbf{o}_{self}$, $\mathbf{o}_{drones}$, $\mathbf{o}_{static}$, and $\mathbf{o}_{dynamic}$. Each part is individually encoded through separate MLPs to produce embeddings of identical dimensions. We use multi-head self-attention on these embeddings to obtain respective features. To better capture the relationships between the drone and other environmental entities, these features are processed through a multi-head cross-attention module, with self-feature as the query and other features as the key and value. The final feature $\mathbf{f}$ is a concatenation of the self-feature and the output of the multi-head cross-attention.
V-C2 Actor&Critic Network
@@
Following the observation encoder, we use a Gaussian distribution model $\pi_{\theta}=\mathcal{N}(\bm{\mu}_{\theta},\bm{\sigma}_{\theta})$ to parameterize actions in the actor network, and an MLP $V_{\phi}(\mathbf{f})$ to estimate the value for the current state in the critic network.
@@ VI Experiments
We compare our approach with three state-of-the-art baselines that meet some objectives: Swarm-Formation [16], R-Mader [9], and Swarm-RL [34]. The first two are planning-based algorithms, and the third one is RL-based algorithm. Since all baselines fail to meet all four objectives in our task, we make the following adaptations:
  • Swarm-Formation [16] is designed to handle formation maintenance around static obstacles. We integrate a trajectory prediction module to facilitate dynamic obstacle avoidance. Upon detecting a dynamic obstacle, the module captures two sequential frames of the obstacle and fits a parabolic curve to estimate its trajectory. This estimated parabolic curve is then treated as a static obstacle and sent to the planner.
  • R-Mader [9] focuses on avoiding both static and dynamic obstacles for multiple drones without considering formation. To incorporate formation, we set the desired formation as the target positions of the drones. For dynamic obstacle avoidance, R-Mader requires knowledge of the obstacles’ trajectories. Therefore, we fit parabolic curves to the obstacles and provide the curve parameters to R-Mader.
  • Swarm-RL [34] focuses on collision avoidance for static obstacles and inter-drone collisions but does not handle dynamic obstacles or formation maintenance. To address this, we incorporate the relative position and velocity of dynamic obstacles into its observation. Additionally, we introduce a velocity tracking reward to aid in training. As with R-Mader, we set the desired formation as the final positions of each drone.
VI-A Simulation Results
We choose OmniDrones [37] as our simulator because of its high-speed GPU simulation and realistic environment dynamics for drones.
We consider a challenging task: 3 drones fly as a near-equilateral triangle while avoiding static and dynamic obstacles scattered within the space. We obtain the policy through our proposed 2-stage RL training pipeline. During the first training stage, each trial requires 100 million environment steps. During the second training stage, we perform 3-period curriculum learning, where we train the RL policy in scenarios with no obstacle, 10 static obstacles only, and mixed obstacles with 10 static and 2 dynamic obstacles, sequentially. Each period takes 15 million, 150 million, and 150 million environment steps, respectively. Swarm-RL takes 1 billion environment steps to converge, about 3x of our method.
VI-A1 Evaluation Metrics
@@
  • Formation Maintenance (FM). In successful episodes, we compute the unnormalized Laplacian distance $e_{shape}$ between the target formation and the actual swarm configuration averaged over the episode. Unlike the normalized metric used during training, this measurement is rotation-invariant but sensitive to size. Hence, a smaller unnormalized Laplacian distance indicates a more desirable formation in terms of both shape and size.
Note that we do not quantify the performance of action smoothing, since this term facilitates sim2real transfer, which will be verified in real-world experiments. In the following section, we report the CFR and FM of planning-based methods averaged over 25 different trials, as they do not support parallel evaluation, and RL-based methods over 100 trials. The positions of all obstacles are randomly generated in each trial.
VI-A2 Main Results
We test all methods in the mixed obstacle scenario with 2 balls and 10 columns and report the performance in Tab. I. We further test the zero-shot transfer ability of all methods by varying the number of obstacles. We gradually increase the number of columns from 5 to 20 under static obstacle scenario, and the number of balls from 1 to 5 under dynamic obstacle scenario. The results are shown in Fig. 3 and Fig. 4, respectively.
@@ VI-A2 Main Results
TABLE I: Performance of all methods under mixed obstacle scenario (2 balls + 10 columns) in simulation.
@@
(a) Static obstacle scenario with 4 columns. The black boxes represent static obstacles.
(b) Dynamic obstacle scenario with 1 ball. The red circles represent the trajectory of dynamic obstacles.
Figure 5: Behavior visualization.
@@
Our method outperforms all baseline methods regarding both CFR and FM in the mixed obstacle scenario, and demonstrates comparable, if not superior, performance to the best baseline in extremely confined environments with 20 columns or 5 balls. The results demonstrate our method’s strong capability to maintain effective formation while evading various obstacles, underscoring its ability to accomplish multi-objective tasks. Also, as the number of obstacles increases, our method shows only a slight degradation in both metrics, demonstrating its strong scalability and generalization capability, which we attribute to the attention mechanism employed in the network structure.
R-Mader, which inherently incorporates dynamic obstacle avoidance, performs better than Swarm-Formation in dynamic scenarios but still falls short of our method. Like Swarm-Formation, R-Mader struggles to navigate through dense distributions of fine-grained obstacles, leading to a low CFR in static scenarios. Moreover, R-Mader is not designed for formation maintenance. As a result, the close proximity of drones in a dense formation makes it difficult to find a collision-free solution.
VI-A3 Behavior Visualization
In Fig. 5(a), we visualize the behaviors of different methods in the same scenario with 4 columns. Our method achieves obstacle avoidance by adjusting the centroid of the formation and slightly twisting its shape, demonstrating its flexibility in formation maintenance. Swarm-Formation and R-Mader fail to find paths around the fine-grained obstacles. Therefore, the green drone in Swarm-Formation and the orange drone in R-Mader remain stationary when other drones pass through. Swarm-RL fails to maintain formation during obstacle avoidance.
Fig. 5(b) visualizes behaviors in scenario with 1 ball. Except for Swarm-Formation, all methods successfully avoid the ball. Our method maintains formation by having all drones decelerate upon encountering the ball and then resume flight after it drops. In R-Mader, the orange drone decelerates and falls behind the other drones, thereby disrupting the formation. In Swarm-RL, the orange drone circles around the ball and rejoins the other drones after avoiding the obstacle.
VI-B Analysis of Reward Scalarization
Fig. 6 (a) illustrates the satisfaction rates of policies trained with randomly sampled weight vectors w. We observe that weight vectors with similar satisfaction rates can yield diverse behaviors. Fig. 6 (b) depicts the trajectories of 3 drones dodging a dynamic obstacle, guided by policies with similar satisfaction rates. When encountering a dynamic obstacle, the first policy and the third policy decelerate the drones in different magnitudes, whereas the second policy choose to let the drones ascend and accelerate. This suggests that our method holds strong flexibility and policy diversity, allowing humans to choose behaviors that best align with their preferences. In practice, we select the weight vector with the highest satisfaction rate.
VI-C Ablation Studies
@@
In this section, we perform ablation studies on the attention-based observation encoder and curriculum learning used in the second stage. All experiments are conducted in a mixed obstacle scenario with 10 columns and 2 balls. For each objective, we define success criteria and record the proportion of episodes that meet these criteria. Additionally, we calculate the Satisfaction Rate, which represents the proportion of episodes that satisfy all criteria simultaneously.
VI-C1 Attention-based Observation Encoder
We compare three variants: (1) Ours w/o CA, where the cross-attention (CA) module is replaced with an MLP, (2) Ours w/o SA, where the self-attention (SA) module is replaced with an MLP, and (3) MLP, where both the cross-attention and self-attention modules are substituted with an MLP.
@@ VI-C2 Curriculum Learning
VI-D Real World Deployment
We deploy our RL policy and the planning-based baselines on Crazyflie 2.0. Due to the limited processing power of Crazyflie’s onboard processors, we execute the RL policy and planning baselines on a computer, and send commands to the drones via Crazyradio. We acquire the ground truth positions of the drones and obstacles through a motion capture system and estimate their linear and angular velocities by calculating the differences in position and rotation between consecutive frames. For the RL policy, CTBR commands are transmitted via radio at 50 Hz, while for planning baselines, trajectory waypoints are sent through position commands at 10 Hz. Since Swarm-RL generates thrust commands at 100Hz that exceeds the communication restraints imposed by the radio, we do not deploy it on the real drone.
VI-D1 Sim2Real Transfer
For better Sim2Real transfer, we add domain randomization and CTBR clip during training. More specifically, we randomize the initial position of each drones within 0.05m of the original formation, and randomize the initial rotation within $[-0.1\pi,0.1\pi]$. Also, to avoid extreme actions, we clip the policy output, setting $\omega$ within $[-\pi/4,\pi/4]$ rad/s and $c$ within $[0.4,0.9]$*max_thrust.
VI-D2 Results and Analysis
Swarm-Formation R-Mader Ours
Static (2 columns) 0.5 0.6 0.7
Dynamic (1 ball) 0.3 0.8 0.7
Mixed (2 columns & 1 ball) 0.3 0.5 0.8
TABLE II: CFR of all methods in real world deployment.
In real world deployment, we consider three scenarios, i.e., with 2 static columns, with one manually thrown ball, and with both the columns and the ball. We repeat each experiment ten times and report the CFR of each method in Table II. For full video demonstration, please refer to the supplementary materials or the website.
The results further demonstrate the strong Sim2Real transfer capability of our method. Our method can perform aggressive dodging strategy to safely navigate through all obstacles while maintaining formation, whereas the baseline methods often fail to find a viable path. Swarm-Formation halts when the dynamic obstacle drops before the drone; R-Mader struggles with cluttered static obstacles and fails to keep formation. Through the zero-shot Sim2Real deployment, we validate the efficacy of the action smoothing objective.
@@ VII Conclusion and Future Work
In this paper, we employ multi-agent reinforcement learning to tackle the challenging task of maintaining formation among multiple UAVs while avoiding both static and dynamic obstacles during directed flight. Our approach involves a two-stage training pipeline. In the first stage, we search for a balanced weight of all objectives in a simplified task setting. In the second stage, we apply curriculum learning to navigate the large exploration space, deriving a feasible policy for the more complex task. We introduce an attention-based observation encoder to effectively capture the features of the drones and the environments, ensuring the scalability of the policy. Our approach outperforms the state-of-the-art baselines in terms of both CFR and FM. The effectiveness of our method is further validated through real-world deployment.